# Modelling Multimodal Integration in Human Concept Processing with Vision-Language Models

### v2

@@ Abstract
Text representations from language models have proven remarkably predictive of human neural activity involved in language processing, with the recent transformer-based models outperforming previous architectures in downstream tasks and prediction of brain responses. However, the word representations learnt by language-only models may be limited in that they lack sensory information from other modalities, which several cognitive and neuroscience studies showed to be reflected in human meaning representations. Here, we leverage current pre-trained vision-language models (VLMs) to investigate whether the integration of visuo-linguistic information they operate leads to representations that are more aligned with human brain activity than those obtained by models trained with language-only input. We focus on fMRI responses recorded while participants read concept words in the context of either a full sentence or a picture. Our results reveal that VLM representations correlate more strongly than those by language-only models with activations in brain areas functionally related to language processing. Additionally, we find that transformer-based vision-language encoders—e.g., LXMERT and VisualBERT—yield more brain-aligned representations than generative VLMs, whose autoregressive abilities do not seem to provide an advantage when modelling single words. Finally, our ablation analyses suggest that the high brain alignment achieved by some of the VLMs we evaluate results from semantic information acquired specifically during multimodal pretraining as opposed to being already encoded in their unimodal modules. Altogether, our findings indicate an advantage of multimodal models in predicting human brain activations, which reveals that modelling language and vision integration has the potential to capture the multimodal nature of human concept representations.
footnotetext: *   Equal contributionfootnotetext: †  Corresponding author
@@ 1 Introduction
How does the brain represent semantic knowledge? Whereas many computational models of language still build on the idea that meaning can be extracted from text corpora (Piantadosi and Hill, 2022), increasing evidence from cognitive science and neuroscience suggests that human semantic representations are in fact grounded in sensory experiences. Theories of embodied simulation argue that comprehending and producing words involves both sensory-motor and linguistic information (Louwerse, 2011; Barsalou, 1999; Harnad, 1990; Bergen, 2012). This view is supported by developmental studies on word meaning acquisition (Landau, Smith, and Jones, 1998; Bornstein et al., 2004) as well as behavioural and neuroimaging experiments highlighting the role of perceptual attributes in processing word meanings (Barsalou, 2008). The need to ground meaning representations in the physical world has also been emphasized by researchers in natural language processing as an important step towards better language understanding in machines (Bender and Koller, 2020; Bisk et al., 2020).
In human concept processing, visual and linguistic stimuli have been shown to mutually influence each other (Boutonnet and Lupyan, 2015; Noorman, Neville, and Simanova, 2018; Huettig, Rommers, and Meyer, 2011) and also evoke shared neural activity patterns, thought to reflect some form of abstracted, modality-independent conceptual representations (Devereux et al., 2013; Simanova et al., 2014; Popham et al., 2021; Kaup et al., 2024). Although the existence of such modality-independent activity patterns does not necessarily tell us much about their representational content, recent evidence indicates that they indeed encode both linguistic and visual information (Dirani and Pylkkänen, 2024). For example, Tang et al. (2024) showed that brain activity in visual processing areas can be successfully predicted by computational models trained to predict the same kind of responses during language processing; the effect is even stronger when such models integrate stimulus features from both modalities. Such analyses, comparing between brain activity patterns and representations generated by deep neural network (DNN) models, offer a promising way to shed further light on the nature of multimodal integration in the brain.
In this paper, we investigate the ability of vision-language models (VLMs) to capture multimodal aspects of human semantic processing. From the perspective of technical applications, there is no doubt that VLMs can perform tasks, such as visual question answering and image captioning, that are simply impossible for language-only models. This opens up the intriguing question of whether VLMs also learn text representations that model human language processing more accurately.
Regarding human behavioural responses, Pezzelle, Takmaz, and Fernández (2021) found promising evidence that VLM embeddings correlate with human similarity judgments more strongly than embeddings from language-only models when the judgments concern highly concrete concepts. As for brain responses, some recent works have explored the brain-modelling potential of VLMs using a variety of stimuli—including pictures Conwell et al. (2023a); Doerig et al. (2022); Wang et al. (2023), audio stories Tang et al. (2023), movies Tang et al. (2023); Dong and Toneva (2023); Subramaniam et al. (2024), and combinations of pictures and textual stimuli Oota et al. (2022); Nikolaus et al. (2024)—and a variety of experimental procedures, such as brain encoding Dong and Toneva (2023); Subramaniam et al. (2024); Oota et al. (2022), cross-encoding Tang et al. (2023) and cross-decoding Nikolaus et al. (2024). However, most of these studies focus on brain responses elicited in complex multimodal settings (e.g., watching movies or listening to long stories), as opposed to linguistic processing of single concepts.
Closer to our focus, Oota et al. (2022) used VLMs as the basis for encoding models of functional magnetic resonance imaging (fMRI) responses to concepts presented through written words and images simultaneously. We extend this work by additionally analysing fMRI responses in a purely linguistic condition and evaluating additional VLM architectures, while using a simpler method for measuring representational alignment. More concretely, we extract representations of concept words with a set of pre-trained VLMs, and apply Representational Similarity Analysis (RSA, Kriegeskorte, Mur, and Bandettini, 2008) to compute their alignment with human fMRI responses recorded while participants are viewing written concept words, either accompanied by pictures (picture condition) or embedded in sentences (sentence condition); see Figure 1 for illustration.
Since vision-language models largely differ in architecture and specific strategies employed to achieve multimodal integration, we identify three main families—contrastive VLMs, vision-language encoders, and generative VLMs—and include representatives from each in our experiments. In addition, we evaluate a set of language-only architectures that roughly match the language model component in the VLMs we consider.
In summary, we investigate two key research questions:
  1. (1)
    Are representations by current pretrained VLMs better models of brain activity involved in concept word processing than text-only language models? As mentioned above, human concept representations seem to reflect knowledge from different modalities. Since VLMs are trained to align input from visual and linguistic streams, we hypothesise they will exhibit an advantage over text-only language models in modelling brain activity during concept processing.
  2. (2)
    How does the brain alignment of different VLM families vary across conditions and brain networks? The experimental conditions we consider differ fundamentally in that the target concept words are presented within either a sentential or a visual context. The presence vs. absence of a sentential context is known to affect the way human brains respond to language Xu et al. (2005); similarly, the ability to process longer contexts is known to result in stronger capabilities at downstream tasks in DNN models. We conduct an extensive comparison of VLM architectures and families, considering three brain networks for each condition; this allows us to understand which specific aspects of brain language processing are better captured by each model.
Altogether, the present work makes the following contributions:
  • It provides, for the first time, a systematic and exhaustive investigation of the brain alignment achieved by transformer-based pretrained VLMs belonging to different model families;
  • It shows evidence that the highest brain alignment is consistently achieved by one of the VLMs (and not a language-only model), although not the same architecture across all conditions and brain networks;
  • It reveals that vision-language encoders, such as LXMERT Tan and Bansal (2019) and VisualBERT Li et al. (2019), tend to exhibit higher brain alignment than the more recent generative VLMs;
  • It demonstrates that the superior brain alignment achieved by vision-language encoders stems from learning novel semantic information and cannot simply be attributed to receiving additional visual input or to semantic information already present in their language-only backbone.
We open-source code and materials to reproduce our experiments at the following public GitHub repository: https://github.com/dmg-illc/concept-processing-brain-vlms.
2 Background
2.1 Use of deep neural networks to model human brain activity
DNN models trained on large amounts of data with task-optimized (or, alternatively, self-supervised) representation-learning algorithms currently underlie virtually all state-of-the-art AI systems. Such models have enabled unprecedented advances in the past decade across the fields of computer vision and natural language processing, as well as applications integrating the two Mitchell et al. (2019). Meanwhile, computational cognitive (neuro)science has made impressive progress using DNNs as model systems of human behaviour and its neural substrates. Traditional experimental approaches often rely on very specific theory-driven hypotheses, with accordingly designed controlled stimuli and contrasting conditions. On the other hand, DNNs provide a complementary approach to studying human information processing, by generating rich representational spaces to encode more complex naturalistic stimuli. In a relatively new line of studies, DNNs are employed as stimulus-computable models, whose internal representations and output behaviour are directly compared to human neural and behavioural responses measured when fed with similar input (Kriegeskorte, Mur, and Bandettini, 2008; Ma and Peters, 2020; Doerig et al., 2023).
Internal representations from DNNs have proven remarkably predictive of human brain activity in both visual and linguistic processing. In the domain of vision, early studies using convolutional neural networks (CNNs) optimised for object recognition found that representations from such models successfully predict fMRI (Güçlü and Van Gerven, 2015; Eickenberg et al., 2017), magnetoencephalography (MEG; Seeliger et al., 2018) and neural spiking (Yamins et al., 2014) responses to natural images. More recent work comparing CNNs to the larger vision-transformer architectures has shown that performance at predicting visual cortex activity is in fact comparable between models with more human-like architectural biases (e.g., convolutional processing mechanisms, as in CNNs) and models with more human-like training objectives (e.g., no direct category supervision, as in vision transformers), when training data is controlled for (Conwell et al., 2023b).
Regarding models of language, early work pioneered the use of corpus-derived and neural network-based static word embeddings as predictive models of brain activity in linguistic tasks such as word reading/processing (Mitchell et al., 2008), sentence reading (Pereira et al., 2018) and narrative understanding (Wehbe et al., 2014). However, a huge improvement in this domain has been achieved by employing internal states from LSTM- and, later, transformer-based language models (e.g., Jain and Huth 2018; Toneva and Wehbe 2019; see Karamolegkou, Abdou, and Søgaard 2023 for a review), whose ability to learn contextualised representations is more akin to the human online processing of linguistic stimuli in context. Most prominently, Schrimpf et al. (2021) compared a comprehensive range of neural network-based embeddings—including static GloVe embeddings (Pennington, Socher, and Manning, 2014), LSTMs (Hochreiter and Schmidhuber, 1997), BERT-based masked language models (Devlin et al., 2019) and GPT-like autoregressive language models (Radford et al., )—in their predictivity of human neural (functional magnetic resonance imaging and electrocorticography) and behavioural (reading time) data and found that generative, autoregressive architectures were most predictive of human cognitive signals in language processing.
The wide applicability of pre-trained DNN hidden state representations and derived model output measures highlights one of the major strengths of this class of models compared to other cognitive modelling paradigms: Models trained to capture a rich variety of phenomena are more likely to be useful for multiple down-stream tasks and also to overlap in interesting ways with similar general-purpose representations involved in human language processing (Antonello and Huth, 2024).
2.2 Early multimodal models of meaning representations
Despite successes in applying DNNs to neural and cognitive modelling across the visual and linguistic domains, most studies described above concern unimodal DNNs, encoding either visual or textual input but not both. However, earlier approaches in computational semantics have repeatedly found that models of lexical meaning that combine visual and textual information are better aligned with human word representations than their unimodal counterparts. Many studies in this line of work combined modalities by concatenating textual and perceptual vectors, often drawing on separate sources of data for building concept representations in either domain (Bruni et al., 2012; Kiela and Bottou, 2014; Kiela, Verő, and Clark, 2016; Derby et al., 2018; Davis et al., 2019); in parallel, there have been efforts to construct unified multimodal embeddings aimed at representing the latent information shared across textual and perceptual modalities (Lazaridou, Pham, and Baroni, 2015; Zablocki et al., 2018; Silberer and Lapata, 2012, 2014).
Overall, such multimodal distributional semantic models were shown to outperform text-only approaches at tasks requiring visual knowledge, as well as more general semantic benchmarks measuring, for example, correlations to human word similarity or relatedness judgments (e.g., WordSim-353, Finkelstein et al. 2001; SimLex999, Hill, Reichart, and Korhonen 2015; MEN, Bruni et al. 2012). Moreover, concatenated visio-linguistic embeddings also showed an advantage over text- and vision-only baselines in decoding brain activity related to concrete noun representations (Davis et al., 2019). More recent studies have evaluated representations extracted from fully end-to-end models trained to map between images and corresponding text captions and found that such models show improved performance over text models on sentence representation tasks (Kiela et al., 2018) as well as improved alignment to human word similarity and relatedness judgments (Merkx, Frank, and Ernestus, 2022). These findings beg the question of whether such emergent cross-/multimodal representations in end-to-end models are also well-aligned with brain activity involved in human concept processing.
2.3 Multimodal semantic representations in the era of vision-language transformers
In recent years, a powerful class of transformer-based VLMs have achieved outstanding performance on a variety of downstream AI tasks. After being pre-trained on matched visual and textual data (e.g., image-caption pairs), these models turn out particularly capable at tasks requiring integrating information from textual and visual inputs, such as visual question answering, image captioning, and visual dialogue (see Suglia, Konstas, and Lemon 2024 for a review). In addition to such broad applications, VLMs have also been evaluated on specific aspects of human-like semantic and situational understanding, such as compositionality (Thrush et al., 2022a), goal-based action selection (Pezzelle et al., 2020) and embodied simulation (Jones and Trott, 2024), as well as their capabilities to deal with particular linguistic phenomena, such as gradable adjectives (Pezzelle and Fernández, 2019) and other constructions (Parcalabescu et al., 2022a; Chen, Fernández, and Pezzelle, 2023). These studies indicate that, despite their success in downstream applications, VLMs often lag behind humans on such targeted linguistic evaluations, opening the question of how closely their semantic representations are truly aligned with human ones.
Beyond their applications in AI, VLMs have also been evaluated in their ability to capture information relevant for modelling human brain responses. A few works on visual processing have demonstrated the advantages of VLMs over vision-only models in matching human brain responses. For instance, Wang et al. (2023) showed that image embeddings from CLIP (a model trained to match embeddings of images and corresponding text captions in a joint space; Radford et al., 2021) are significantly more predictive of high-level visual cortex activity in picture viewing than models trained on simpler image-label pairs or text only.
Another line of work has used VLMs to model brain responses to more complex multimodal stimuli, such as movies, stories, and combinations of images and text, offering mixed evidence regarding multimodal models’ ability to learn information beyond that learnt by unimodal models. For example, Nikolaus et al. (2024) focussed on cross-decoding fMRI responses to pictures and their textual descriptions and reported that multimodal representations do not increase decoding performance compared to unimodal (vision-only and language-only) representations. Similarly, Dong and Toneva (2023) used a multimodal transformer (processing video frames, text, and audio) to model fMRI recordings from participants watching a popular TV show and found that vision-audio representations were more predictive of brain activity in language regions than audio-only and vision-only representations; however, the brain predictive power of the multimodal representations did not go beyond the combination of that of unimodal representations. This partially contrasts with the findings by Tang et al. (2023), who used representations from VLMs to train encoding models generalising between fMRI responses to movies and audio stories about the same contents, and showed that this encoding performance was significantly higher than that by unimodal representations (vision-only and language-only). Finally, Subramaniam et al. (2024) used representations from VLMs to predict stereo-electroencephalography (sEEG) responses to movies, and identified sites of multimodal brain integration by checking in which areas the VLM representations were more predictive than language-only, vision-only, and linearly-combined vision and language representations.
@@ 2.3 Multimodal semantic representations in the era of vision-language transformers
Moving to single-word processing, Pezzelle, Takmaz, and Fernández (2021) evaluated word representations extracted from several VLM architectures against human semantic judgments, including the word similarity and relatedness benchmarks mentioned in Section 2.2. VLM word representations aligned better with human judgments than representations from text-only models, although only on concrete word pairs. Furthermore, Zhuang, Fedorenko, and Andreas (2024) evaluated a set of language-only, vision-only, and vision-language architectures on several measures of word learning over model training, and found that visual information is beneficial only when small amounts of training data are available. However, none of these works involved brain responses.
Taken together, these works suggest that VLMs are capable—although not always more capable than unimodal models—to predict brain responses to multimodal stimuli and capture some cognitively relevant aspects of single-word representations. A natural next question is whether VLMs are better than language-only models at predicting brain responses from isolated concept words. This issue was partially addressed by Oota et al. (2022), who predicted fMRI responses to single concepts—presented through both images and text—with representations by unimodal (language-only and vision-only) models and VLMs against fMRI responses and found VisualBERT Li et al. (2019) to be more brain-predictive than the other architectures. We extend this work by including fMRI responses to purely textual stimuli and evaluating representations from additional, more recent VLMs.

### v3

# Vision-Language Models Align with Human Neural Representations in Concept Processing

@@ Abstract
Recent studies suggest that transformer-based vision-language models (VLMs) capture the multimodality of concept processing in the human brain. However, a systematic evaluation exploring different types of VLM architectures and the role played by visual and textual context is still lacking. Here, we analyse multiple VLMs employing different strategies to integrate visual and textual modalities, along with language-only counterparts. We measure the alignment between concept representations by models and existing (fMRI) brain responses to concept words presented in two experimental conditions, where either visual (pictures) or textual (sentences) context is provided. Our results reveal that VLMs outperform the language-only counterparts in both experimental conditions. However, controlled ablation studies show that only for some VLMs, such as LXMERT and IDEFICS2, brain alignment stems from genuinely learning more human-like concepts during pretraining, while others are highly sensitive to the context provided at inference. Additionally, we find that vision-language encoders are more brain-aligned than more recent, generative VLMs. Altogether, our study shows that VLMs align with human neural representations in concept processing, while highlighting differences among architectures. We open-source code and materials to reproduce our experiments at: https://github.com/dmg-illc/vl-concept-processing.
Vision-Language Models Align with
Human Neural Representations in Concept Processing
Anna Bavaresco, Marianne de Heer Kloots, Sandro Pezzelle, Raquel Fernández Institute for Logic, Language and Computation University of Amsterdam {a.bavaresco, m.l.s.deheerkloots, s.pezzelle, raquel.fernandez}@uva.nl
@@ 1 Introduction
Recent attempts to augment language-only models with vision have resulted in a multitude of vision-language models (VLMs), integrating modalities with different strategies based on the task at hand. While the practical implications of multimodality are evident—it allows language-only models to ‘see’ and perform novel vision-language tasks—, its theoretical implications from the point of view of human language modelling are not yet fully understood.
Insights from cognitive and neuroscientific work suggest that human semantic representations are deeply grounded in multimodal sensory experiences Louwerse (2011); Barsalou (1999); Harnad (1990); Bergen (2012), and that visual and linguistic stimuli evoke shared neural activity patterns (Devereux et al., 2013; Simanova et al., 2014; Popham et al., 2021; Kaup et al., 2024). These findings motivate the hypothesis that semantic representations learnt by VLMs approximate the human ones better than representations by unimodal (both language-only and vision-only) models.
However, existing works testing this empirically point to diverging conclusions, with some studies documenting advantages of multimodality (e.g., Oota et al., 2022; Zhuang et al., 2024a), others suggesting that multimodal fine-tuning ‘harms’ the human-alignment properties originally acquired by language models (e.g., Bavaresco and Fernández, 2025), and more nuanced contributions indicating that the advantages of multimodality are circumscribed (e.g., Pezzelle et al., 2021; Dong and Toneva, 2023; Zhuang et al., 2024b; Ryskina et al., 2025). A likely reason behind the difficulties in reconciling the findings from these studies lies in their addressing similar research questions with different methods.
Among others, two under-investigated experimental factors that may strongly impact measures of human-likeness computed on VLM semantic representations concern the type and amount of context provided, and the differences between VLM architectures. More concretely, many psycholinguistic studies on multimodality in concept processing analysed human measures (both behavioural and neural) collected by presenting participants with non-contextualised words (e.g., Pezzelle et al., 2021; Zhuang et al., 2024b; Bavaresco and Fernández, 2025). This may have resulted in limited engagement of multimodal knowledge, ultimately leading to an under-appreciation of VLMs’ brain modelling potential. An additional limitation of previous work is that it mostly focused on the multimodal/unimodal dichotomy, placing little emphasis on the substantial differences existing between multimodal architectures, and their implications for brain modelling.
In this study, we systematically compare ten models, including VLMs and language-only counterparts, by measuring their correlation with a subset of the Pereira dataset Pereira et al. (2018), collecting human fMRI (functional magnetic resonance imaging) responses to 180 concepts. Our experiments address the limitations of previous work in two respects.
First, we explore the role played by the context provided through different modalities; we analyse brain responses to concept words that were presented to participants with either a sentential context (sentence condition, where concept words are presented within sentences) or a visual one (picture condition, where concept words are accompanied by images illustrating their content), as shown in Figure 1. Besides ensuring that multimodal knowledge is engaged, focusing on this dual-context setup allows us to better characterise the brain-aligning properties of VLMs: if VLMs are more brain-aligned than their language-only counterparts only when visual context is provided, their superiority may be simply due to accessing visual information that language-only models cannot ‘see’; on the other hand, if VLMs have an advantage even without receiving an image input, this signals they have truly learnt human-like, multimodal concept representations during pretraining.
Second, we evaluate different VLM families: we indentify three classes of VLMs integrating the visual and textual modalities through different strategies, and test two representative exemplars for each. By comparing their alignment with brain responses, we test whether multimodality leads to a generalised pattern of improvement or is family/architecture-dependent.
Our results reveal that VLMs tend to be more brain-aligned than their language-only counterparts in both context conditions, suggesting that they successfully model human-like concept processing. In addition, we find that vision-language encoders, such as VisualBERT Li et al. (2019) and LXMERT Tan and Bansal (2019), model brain responses better than the other VLM types, including the more powerful, generative ones. This means that performance on downstream tasks, where generative VLMs excel, may not go hand in hand with human-like concept processing. More broadly, our findings contribute to the ongoing debate around whether multimodality results in more human-like language models or not, showing that it is beneficial for brain alignment in concept processing.
2 Related Work
Figure 1: Overview of the experimental setup in the sentence (top) and picture (bottom) condition. Models are fed with the same stimuli participants saw in the fMRI scanner, i.e., concept words appearing in six contexts (provided by either sentences or pictures). Note that contexts are intended to highlight the same word meaning, but may describe different situations (sentences are not image captions). Model representations and brain responses averaged across the six contexts are then used to derive representational dissimilarity matrices (RDMs), storing pairwise cosine distances. Finally, the Spearman correlation between these RDMs provides a measure for model–brain alignment. Best viewed in colour.
2.1 Multimodal models in cognitive modelling
Several recent studies have used transformer-based multimodal models in brain encoding/decoding experiments or to model behavioural data. Here, we review the closest to our focus, i.e., those analysing responses to concepts as opposed to more complex stimuli (e.g., videos).
@@ 2.1 Multimodal models in cognitive modelling
Pezzelle et al. (2021) evaluated word representations extracted from several VLM architectures against human semantic judgments, including word similarity and relatedness benchmarks. Their findings reveal that VLM word representations align better with human judgments than representations from text-only models, although only on concrete word pairs. Similarly, Zhuang et al. (2024a) and Zhuang et al. (2024b) found visual grounding to result in improved and more efficient word learning, especially in low-data regimes.
Considering studies analysing brain activations, Oota et al. (2022) used several unimodal (language-only and vision-only) models and VLMs to predict a subset of the Pereira dataset (only the picture condition), and found a VLM (VisualBERT, Li et al., 2019) to be more brain-predictive than the other architectures. Crucially, in this setup, multimodal representations are the only ones receiving the same input shown to the human participants (i.e., image+word). Therefore, it is difficult to determine if their advantage is due to capturing something fundamental about concept processing, or simply to having access to more information than unimodal models.
Bavaresco and Fernández (2025) partially addressed this limitation by comparing multimodal and language-only models in a controlled setup, where they were all fed with text-only inputs and used to model fMRI responses collected while participants viewed isolated nouns. Their results, surprisingly, reveal an advantage of language-only models, even on a subset of more concrete nouns. As it is known that the presence of linguistic context influences both brain responses to concepts Xu et al. (2005); Deniz et al. (2023) and mental simulations of their content Zwaan (2014), a likely explanation for their findings is that the brain responses they analysed did not reflect a detectable engagement of multimodal knowledge.
Finally, Ryskina et al. (2025) analysed, again, the Pereira dataset and predicted fMRI responses using both VLMs and large language-only models (LLMs). Their main goal was to identify concept-sensitive brain regions and check whether their neural activations can be successfully modelled with VLMs and LLMs. Systematic comparisons of VLM families and analyses of the role played by context were, therefore, beyond the scope of their study.
We complement these works by analysing brain responses to concepts presented within different context conditions, including a picture condition, as analysed by Oota et al. (2022), and a sentence condition. This allows us to determine if VLMs’ alignment with human brain responses is modulated by the context or consistent across visual and sentential contexts.
2.2 VLM families
Transformer-based vision-language models (VLMs) can be divided into three categories: contrastive models, vision-language encoders, and generative VLMs. Contrastive models, such as CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and LiT Zhai et al. (2022), encode images and text separately, with two dedicated transformer-based modules. These modules are pre-trained with a contrastive loss, which maximises the similarity between image and text embeddings of matching image-text pairs.
Vision-language encoders are characterised by a specific module that uses attention mechanisms to learn relations between visual features extracted with an object detector and text embeddings. In some architectures, such as VisualBERT Li et al. (2019), language processing and multimodal integration are performed by the same BERT-based Devlin et al. (2019) module; in other architectures, e.g., LXMERT Tan and Bansal (2019), text is encoded in a dedicated transformer module before being passed to the cross-modal module.
Lastly, generative VLMs consist of a pretrained LLM, a pretrained vision encoder, and an ‘adaptor’ (or ‘projector’), i.e., a shallow module which learns a mapping between the space of the image tokens and that of the language tokens. Examples of these architectures are LLaVA-NeXT Liu et al. (2024), IDEFICS2 Laurençon et al. (2024), Qwen2.5-VL Bai et al. (2025), and Molmo Deitke et al. (2024). [^1] Here, we do not review proprietary models as they do not allow extracting layer-wise representations, which are fundamental for assessing brain alignment.
The studies reviewed above (Section 2.1) used exemplars from at most two of the VLM families, but a comprehensive comparison remains elusive. To address this gap, we experiment with six different VLMs, including representatives from each family.